Pecos Manufacturing has just issued a 15​-year, 12​% coupon interest​ rate, ​$1000​-par bond that pays interest annually.The required return is currently 13​%, and the company is certain it will remain at 13​% until the bond matures in 15 years.

a.Assuming that the required return does remain at 13​% until​ maturity, find the value of the bond with​ (1) 15 ​years, (2) 12​ years, (3) 9​ years, (4) 6​ years, (5) 3​ years, (6) 1 year to maturity.

b.All else remaining the​ same, when the required return differs from the coupon interest rate and is assumed to be constant to​ maturity, what happens to the bond value as time moves toward​ maturity?
Answer #1

a]

Price of a bond is the present value of its cash flows. The cash flows are the coupon payments and the face value receivable on maturity

Price of bond is calculated using PV function in Excel :

rate = 13% (YTM of bonds = required return)

nper = Years remaining until maturity (1 coupon payment each year)

pmt = annual coupon payment = face value * coupon rate

fv = face value receivable on maturity

The price of bond at different years to maturity is calculated as below :

А в Years to Maturity Price 15 $935.38 12 $940.82 9 $948.68 б $960.02 3 $976.39 1 $991.15

Years to 1 Maturity Price 2 15 =PV(13%,A2, 1000*12%, 1000)*-1 3 12 =PV(13%,A3,1000*12%, 1000)*-1 =PV(13%, A4, 1000*12%, 1000)

b]

As the bond towards maturity, the price of the bond converges (moves towards) to its face value